# Vulnerability Webs:

### v2

@@ A.1.1 Variational EM algorithm with MM updates
$u_{ij,kl}(\bm{\alpha},\bm{\beta})=u(\bm{x}_{i},\bm{x}_{j},z_{ik}=z_{jl}=1,\bm{z};\bm{\alpha},\bm{\beta}).$
Note that $g_{ii}\coloneqq 0$ for all $i$ by definition. The covariate information $\bm{x}$ encodes in our setting categorical information on the level of repositories (maturity, popularity, and size). We, therefore, introduce for the pair of repositories $(i,j)$ the covariate vector $\bm{x}_{ij}=(x_{ij,1},\ldots,x_{ij,p})$, where the $k$th entry is a binary indicator whether repositories $i$ and $j$ match on the $k$th covariate. The variables $x_{p,ij}\coloneqq\bm{1}\{x_{ip}=x_{jp}\}$ are, therefore, indicators for homophily. Generalizations are possible, as long as we maintain the discrete nature of the covariates $\bm{x}$. Including continuous covariates is possible, but this would result in a significant slowdown of this algorithm. Observing (A.1.1), we can drop the dependence on the particular pair of repositories and write $\pi_{kl}(g_{ij},\bm{x}_{ij})=\pi_{ij,kl}(g_{ij},\bm{x})$.
We maximize (A.1.1) in a blockwise manner,first in terms of $\bm{\eta},\bm{\alpha},$ and $\bm{\beta}$ conditional on $\bm{\Xi}$ and then the other way around. Denoting the values of $\bm{\Xi},\bm{\eta},\bm{\alpha},$ and $\bm{\beta}$ in the $t$th iteration by $\bm{\Xi}^{(s)},\bm{\eta}^{(s)},\bm{\alpha}^{(s)},$ and $\bm{\beta}^{(s)}$, the two steps comprising the Variational Expectation Maximization algorithm are:
  • Step 1: Given $\bm{\Xi}^{(s)}$, find $\bm{\alpha}^{(s+1)}$ and $\bm{\beta}^{(s+1)}$ satisfying:
    $\ell_{B}\left(\bm{g},\bm{x},\bm{\alpha}^{(s+1)},\bm{\beta}^{(s+1)},\bm{\eta}^{(s+1)};\bm{\Xi}^{(s)}\right)\,\geq\,\ell_{B}\left(\bm{g},\bm{x},\bm{\alpha}^{(s)},\bm{\beta}^{(s)},\bm{\eta}^{(s)};\bm{\Xi}^{(s)}\right);$
  • Step 2: Given $\bm{\alpha}^{(s+1)}$ and $\bm{\beta}^{(s+1)}$, find $\bm{\Xi}^{(s)+1}$ satisfying:
    $\ell_{B}\left(\bm{g},\bm{x},\bm{\alpha}^{(s+1)},\bm{\beta}^{(s+1)},\bm{\eta}^{(s+1)};\bm{\Xi}^{(s+1)}\right)\,\geq\,\ell_{B}\left(\bm{g},\bm{x},\bm{\alpha}^{(s+1)},\bm{\beta}^{(s+1)},\bm{\eta}^{(s+1)};\bm{\Xi}^{(s)}\right).$
Step 1:
Taking first order conditions with respect to each parameter implies the following closed-form update rules for $\bm{\eta}$ and $\pi_{kl}^{(s+1)}(d,x_{1},\ldots,x_{p})$:
$\eta_{k}^{(s+1)}:=\frac{1}{n}\sum_{i=1}^{n}\xi_{ik}^{(s)},\quad k=1,\ldots,K,$ (25)
and
$\pi_{kl}^{(s+1)}(d,x_{1},\ldots,x_{p}):=\frac{\sum_{i=1}^{n}\sum_{j\neq i}\xi_{ik}^{(s)}\xi_{jl}^{(s)}\bm{1}\{g_{ij}=d,X_{1,ij}=x_{1},\ldots,X_{p,ij}=x_{p}\}}{\sum_{i=1}^{n}\sum_{j\neq i}\xi_{ik}^{(s)}\xi_{jl}^{(s)}\bm{1}\{X_{1,ij}=x_{1},\ldots,X_{p,ij}=x_{p}\}},$ (26)
for $k,l=1,\ldots,K$ and $d,x_{1},\ldots,x_{p}\in\{0,1\}$, respectively. The probability $\pi_{kl}^{(s+1)}(d,x_{1},\ldots,x_{p})$ is the same for all packages $i$ and $j$ such that $z_{i}=k$ and $z_{j}=l$. Therefore, we Note that we are not interested in finding particular estimates for $\bm{\alpha}$, $\bm{\beta}$ and $\gamma$ when running the Variational EM (VEM) algorithm. We are only estimating the probabilities $\pi_{kl}(g,\bm{x})$ for $g\in\{0,1\}$ and $\bm{x}\in\mathcal{X}$, where $\mathcal{X}$ is the set of all possible outcomes of the categorical covariates used in Step 1. This feature allows us to speed up the computation of several orders of magnitude (Dahbura et al., 2021). Finally, remark that updates (25) and (26) can both be comprehended as functions of $\bm{\Xi}^{(s)}$.
Step 2:
For very large networks, maximizing the lower bound in the second step with respect to $\bm{\Xi}$ for given estimates of $\bm{\eta},\bm{\alpha},$ and $\bm{\beta}$ is even under the variational approximation of (22) cumbersome and impractically slow. We thus borrow an idea from Vu et al. (2013) define a function $M\left(\bm{\Xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)$ that minorizes the lower bound $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi})$ in $\bm{\Xi}^{(s)}$. Therefore,
$M\left(\bm{\Xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)$$\leq$ $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi})\text{ \ \ for all }\bm{\Xi}\in[0,1]^{N\times K}$
$M\left(\bm{\Xi}^{(s)};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)$ $=$ $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi}^{(s)})$
holds with fixed $\bm{\alpha},\bm{\beta},\bm{\eta}$ and $\bm{\Xi}^{(s)}$. Applying these two properties guarantees that maximizing the function $M$ forces the lower bound uphill:
$\begin{array}[]{llllllllll}\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi}^{(s)})&=&M\left(\bm{\Xi}^{(s)};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)&\leq&M\left(\bm{\Xi}^{(s+1)};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)\\ &\leq&\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi}^{(s+1)})\end{array}$
Extending the approach of Vu et al. (2013) to directed networks, the following expression minorizes $\ell_{B}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\bm{\Xi})$
$M\left(\bm{\Xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)$ ::=:= $\sum_{i=1}^{N}\sum_{j\neq i}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\left(\xi_{ik}^{2}\frac{\xi_{jl}^{(s)}}{2\xi_{ik}^{(s)}}+\xi_{jl}^{2}\frac{\xi_{ik}^{(s)}}{2\xi_{jl}^{(s)}}\right)\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})$
$+$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\xi_{ik}\left(\log\eta_{k}^{(s)}-\log\xi_{ik}^{(s)}-\frac{\xi_{ik}}{\xi_{ik}^{(s)}}+1\right).$
In Section B, we detail how sparse matrix multiplication can be exploited to speed up the updates of $\bm{\Xi}^{(s+1)}$.
We run this algorithm until the relative increase of the variational lower bound in below a certain threshold to obtain estimates for $\widehat{\bm{\Xi}}$ and $\widehat{\bm{\eta}}$. We then assign each node to its modal estimated type. Therefore, package $i$ is of type $k$, $\widehat{z}_{ik}=1$, if $\widehat{\bm{\Xi}}_{ik}\geq\widehat{\bm{\Xi}}_{i\ell}$ for all $\ell\neq k$ and for all $i$s.
ONLINE APPENDIX
Appendix B Exploiting sparse matrix operations for updating variational parameters
Next, we detail how the updates of $\mathbf{\xi}$ in Step 1 can be carried out in a scalable manner based on sparse matrix operations. Equation (A.1.1) can be written as a quadratic programming problem in $\mathbf{\xi}$ with the block constraints that $\sum_{k=1}^{K}\xi_{ik}=1$ needs to hold for all $i=1,\ldots,N$. We rearrange the first part of (A.1.1) as follows:
$\sum_{i=1}^{N}\sum_{j\neq i}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\left(\xi_{ik}^{2}\,\frac{\xi_{jl}^{(s)}}{2\xi_{ik}^{(s)}}+\xi_{jl}^{2}\,\frac{\xi_{ik}^{(s)}}{2\xi_{jl}^{(s)}}\right)\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})$
$=\sum_{i=1}^{N}\sum_{k=1}^{K}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{ik}^{2}\,\frac{\xi_{jl}^{(s)}}{2\xi_{ik}^{(s)}}\left(\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})+\log\,\pi_{lk}^{(s)}(g_{ji},\bm{x}_{ji})\right)$
$=\sum_{i=1}^{N}\sum_{k=1}^{K}\frac{\Omega_{ik}(\bm{g},\bm{x},\bm{\Xi}^{(s)})}{2\xi_{ik}^{(s)}}\,\xi_{ik}^{2}$
@@ Appendix B Exploiting sparse matrix operations for updating variational parameters
$\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)~{}\coloneqq~{}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left(\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})+\log\,\pi_{lk}^{(s)}(g_{ji},\bm{x}_{ji})\right).$ (28)
This yields for (A.1.1)
$M\left(\bm{\Xi};\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(s)}\right)$ $=\sum_{i=1}^{N}\sum_{k=1}^{K}\left(\frac{\Omega_{ik}^{(s)}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)}{2\,\xi_{ik}^{(s)}}-\frac{1}{\xi_{ik}^{(s)}}\right)\xi_{ik}^{2}+\left(\log\eta_{k}^{(s)}-\log\xi_{ik}^{(s)}+1\right)\xi_{ik}$
$=\sum_{i=1}^{N}\sum_{k=1}^{K}A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)\,\xi_{ik}^{2}+B_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)\,\xi_{ik}$
where
$A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)\coloneqq\frac{\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)}{2\,\xi_{ik}^{(s)}}-\frac{1}{\xi_{ik}^{(s)}}$
is the quadratic term and
$B_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)\coloneqq\log\eta_{k}^{(s)}-\log\xi_{ik}^{(s)}+1$
the linear term of the quadratic problem.
To update the estimate of $\bm{\Xi}$, we need to evaluate $A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ and $B_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ for $i=1,\ldots,N$ and $k=1,\ldots,K$, which, when done naively, is of complexity $O(N^{2}\,K^{2})$. Computing $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ for $A_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ is the computational bottleneck driving these complexities. It is thus prohibitively difficult for a large number of population members and communities in the network. Note that in our application, this problem is exasperated as we have more than 35,000 nodes and our theoretical result on the adequacy of the estimation procedure assumes that $K$ grows as a function of $N$. To avoid this issue, we show how $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ can be evaluated through a series of sparse matrix multiplications. Thereby, our implementation is fast and light on memory requirements. The underlying idea is that $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ collapses to a simple form if $\mathbf{g}$ is an empty network with all pairwise covariates set to zero, i.e., $g_{ij}=0$ and $x_{q}=0\,\forall\,i,j=1,\ldots,N$ and $q=1,\ldots,p$. Given that, in reality, these are seldom the observed values, we, consecutively, go through all connections where either $g_{ij}=1$ or $X_{ij,1}=x_{1},\ldots,X_{ij,p}=x_{p}$ for $\sum_{q=1}^{p}\,x_{q}\neq 0$ holds and correct for the resulting error. In formulae, this implies the following:
$\Omega_{ik}(\bm{g},\bm{x},\bm{\Xi}^{(s)})~{}$ $\coloneqq~{}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left(\log\,\pi_{kl}^{(s)}(g_{ij},\bm{x}_{ij})+\log\,\pi_{lk}^{(s)}(g_{ji},\bm{x}_{ji})\right)$
$=\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left(\log\,\pi_{kl}^{(s)}(0,\bm{0})+\log\,\pi_{lk}^{(s)}(0,\bm{0})\right)$
$+\,\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left(g_{ij}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+g_{ji}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$
$+\,\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left((1-g_{ij})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+(1-g_{ji})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$
$=\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)+\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right),$ (29)
with
$\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ $=\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left(g_{ij}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+g_{ji}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$ (30)
$+\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left((1-g_{ij})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+(1-g_{ji})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$
being the error in $\Omega_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ arising from assuming an empty network with all categorical covariates set to zero. Next, we show how both terms, $\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)$ and $\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$, can be computed through matrix operations, completing our scalable algorithmic approach.
We, first, assume an empty network where all categorical covariates set to zero, $g_{ij}=x_{q}=0$ for all $i\neq j$ and $q=1,...,p$ and compute $\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)$ via matrix multiplications. Observe that
$\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)$ $=\sum_{j\neq i}^{N}\sum_{l=1}^{K}\xi_{jl}^{(s)}\left(\log\,\pi_{kl}^{(s)}(0,\bm{0})+\log\,\pi_{lk}^{(s)}(0,\bm{0})\right)$
$=\sum_{l=1}^{K}\sum_{j\neq i}^{N}\xi_{jl}^{(s)}\left(\log\,\pi_{kl}^{(s)}(0,\bm{0})+\log\,\pi_{lk}^{(s)}(0,\bm{0})\right)$
$=\sum_{l=1}^{K}\left(\sum_{j=1}^{N}\xi_{jl}^{(s)}-\xi_{il}^{(s)}\right)\left(\log\,\pi_{kl}^{(s)}(0,\bm{0})+\log\,\pi_{lk}^{(s)}(0,\bm{0})\right)$
$=\sum_{l=1}^{K}\left(\tau^{(s)}_{l}-\xi_{il}^{(s)}\right)\left(\log\,\pi_{kl}^{(s)}(0,\bm{0})+\log\,\pi_{lk}^{(s)}(0,\bm{0})\right)$
holds, where
$\tau^{(s)}_{l}~{}\coloneqq~{}\sum_{j=1}^{N}\xi_{jl}^{(s)}.$
With
$\bm{A_{0}}\left(\bm{\Xi}^{(s)}\right)~{}\coloneqq~{}\begin{bmatrix}\tau^{(s)}_{1}-\xi_{11}^{(s)}&\tau^{(s)}_{2}-\xi_{12}^{(s)}&\dots&\tau^{(s)}_{K}-\xi_{1K}^{(s)}\\ \tau^{(s)}_{1}-\xi_{21}^{(s)}&\tau^{(s)}_{2}-\xi_{22}^{(s)}&\dots&\tau^{(s)}_{K}-\xi_{2K}^{(s)}\\ \vdots&\vdots&\ddots&\vdots\\ \tau^{(s)}(1)-\xi_{n1}^{(s)}&\tau^{(s)}_{2}-\xi_{n2}^{(s)}&\dots&\tau^{(s)}_{K}-\xi_{nK}^{(s)}\end{bmatrix}$
and
$\bm{\Pi_{0}}\left(\bm{\Xi}^{(s)}\right)~{}\coloneqq~{}\begin{bmatrix}\log\,\pi_{11}^{(s)}(0,\bm{0})&\log\,\pi_{12}^{(s)}(0,\bm{0})&\dots&\log\,\pi_{1K}^{(s)}(0,\bm{0})\\ \log\,\pi_{21}^{(s)}(0,\bm{0})&\log\,\pi_{22}^{(s)}(0,\bm{0})&\dots&\log\,\pi_{2K}^{(s)}(0,\bm{0})\\ \vdots&\vdots&\ddots&\vdots\\ \log\,\pi_{K1}^{(s)}(0,\bm{0})&\log\,\pi_{K2}^{(s)}(0,\bm{0})&\dots&\log\,\pi_{KK}^{(s)}(0,\bm{0})\end{bmatrix},$
it follows that
$\bm{A_{0}}\left(\bm{\Xi}^{(s)}\right)\left(\left(\bm{\Pi_{0}}\left(\bm{\Xi}^{(s)}\right)\right)^{\top}+\bm{\Pi_{0}}\left(\bm{\Xi}^{(s)}\right)\right)=\left(\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)\right)_{ik}$ (31)
holds. Put differently, we are able to write $\Omega_{ik}^{(s)}(\bm{0},\bm{0},\bm{\Xi})$ as the $(i,k)$th entry of the result of (31). To compute $\bm{\Pi_{0}}\left(\bm{\Xi}^{(s)}\right)$, we first evaluate the matrix $\pi^{(s)}(1,\bm{0})\in[0,1]^{K\,\times\,K}$ via sparse matrix operations:
$\bm{\pi}^{(s+1)}(1,\bm{0})=\left(\left(\bm{\Xi}^{(s+1)}\right)^{\top}\bm{g}\circ(\bm{J}-\bm{X_{1}})\circ\cdots\circ(\bm{J}-\bm{X_{p}})\ \left(\bm{\Xi}^{(s+1)}\right)\right)\oslash$ (32)
$\left(\left(\bm{\Xi}^{(s+1)}\right)^{\top}(\bm{J}-\bm{X_{1}})\circ\cdots\circ(\bm{J}-\bm{X_{p}})\left(\bm{\Xi}^{(s+1)}\right)\right),$
$\bm{\Pi_{0}}\left(\bm{\Xi}^{(s)}\right)=\log\left(1-\bm{\pi}^{(s+1)}(1,\bm{0})\right)$
and calculate $\Omega_{ik}^{(s)}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)$ for $i=1,\ldots,N$ and $k=1,\ldots,K$.
$\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)~{}$ $=~{}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left(g_{ij}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+g_{ji}\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$
$+\,\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left((1-g_{ij})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x}_{ij})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)+(1-g_{ji})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x}_{ji})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)\right)$
$=\sum_{\bm{x}\in\mathcal{X}}\sum_{j\neq i}^{N}\sum_{l=1}^{K}g_{ij}\,\mathbb{I}(\bm{x}_{ij}=\bm{x})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(1,\bm{x})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)$
$+\,g_{ji}\,\mathbb{I}(\bm{x}_{ji}=\bm{x})\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(1,\bm{x})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)+(1-g_{ij})\,\mathbb{I}(\bm{x}_{ij}=\bm{x})\,\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{kl}^{(s)}(0,\bm{x})}{\pi_{kl}^{(s)}(0,\bm{0})}\right)$
+ $+\,1-g_{ji})\,\mathbb{I}(\bm{x}_{ji}=\bm{x})\,\xi_{jl}^{(s)}\left(\log\,\dfrac{\pi_{lk}^{(s)}(0,\bm{x})}{\pi_{lk}^{(s)}(0,\bm{0})}\right)$
With
@@
$\bm{\Pi}\left(d,\bm{x},\bm{\Xi}^{(s)}\right)\coloneqq\begin{bmatrix}\log\frac{\pi_{11}^{(s)}(d,\bm{x})}{\pi_{11}^{(s)}(d,\bm{0})}&\log\frac{\pi_{12}^{(s)}(d,\bm{x})}{\pi_{12}^{(s)}(d,\bm{0})}&\dots&\log\frac{\pi_{1K}^{(s)}(d,\bm{x})}{\pi_{1K}^{(s)}(d,\bm{0})}\\ \log\frac{\pi_{21}^{(s)}(d,\bm{x})}{\pi_{21}^{(s)}(d,\bm{0})}&\log\frac{\pi_{22}^{(s)}(d,\bm{x})}{\pi_{22}^{(s)}(d,\bm{0})}&\dots&\log\frac{\pi_{2K}^{(s)}(d,\bm{x})}{\pi_{2K}^{(s)}(d,\bm{0})}\\ \vdots&\vdots&\ddots&\vdots\\ \log\frac{\pi_{K1}^{(s)}(d,\bm{x})}{\pi_{K1}^{(s)}(d,\bm{0})}&\log\frac{\pi_{K2}^{(s)}(d,\bm{x})}{\pi_{K2}^{(s)}(d,\bm{0})}&\dots&\log\frac{\pi_{KK}^{(s)}(d,\bm{x})}{\pi_{KK}^{(s)}(d,\bm{0})}\end{bmatrix}$
and the functions $\bm{\Gamma}\colon\{0,1\}\to\mathbb{R}^{n\times n}$ and $\bm{\Lambda}_{s}\colon\{0,1\}\to\mathbb{R}^{n\times n}$ $(s=1,\ldots,p)$ defined as
$\bm{\Gamma}(d)\coloneqq\begin{cases}\bm{G}&(d=1)\\ \bm{J}-\bm{G}&(d=0)\end{cases}$

### v3

@@ Appendix A Exploiting sparse matrix operations for updating variational parameters
$\bm{\Pi_{0}}\left(\bm{\Xi}^{(t)}\right)~{}\coloneqq~{}\begin{bmatrix}\log\,\pi_{11}^{(t)}(0,\bm{0})&\log\,\pi_{12}^{(t)}(0,\bm{0})&\dots&\log\,\pi_{1K}^{(t)}(0,\bm{0})\\ \log\,\pi_{21}^{(t)}(0,\bm{0})&\log\,\pi_{22}^{(t)}(0,\bm{0})&\dots&\log\,\pi_{2K}^{(t)}(0,\bm{0})\\ \vdots&\vdots&\ddots&\vdots\\ \log\,\pi_{K1}^{(t)}(0,\bm{0})&\log\,\pi_{K2}^{(t)}(0,\bm{0})&\dots&\log\,\pi_{KK}^{(t)}(0,\bm{0})\end{bmatrix},$
it follows that
$\bm{A_{0}}\left(\bm{\Xi}^{(t)}\right)\left(\left(\bm{\Pi_{0}}\left(\bm{\Xi}^{(t)}\right)\right)^{\top}+\bm{\Pi_{0}}\left(\bm{\Xi}^{(t)}\right)\right)=\left(\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(t)}\right)\right)_{ik}$ (22)
holds. Put differently, we are able to write $\Omega_{ik}^{(t)}(\bm{0},\bm{0},\bm{\Xi})$ as the $(i,k)$th entry of the result of (22). To compute $\bm{\Pi_{0}}\left(\bm{\Xi}^{(t)}\right)$, we first evaluate the matrix $\pi^{(t)}(1,\bm{0})\in[0,1]^{K\,\times\,K}$ via sparse matrix operations:
$\bm{\pi}^{(t+1)}(1,\bm{0})=\left(\left(\bm{\Xi}^{(t+1)}\right)^{\top}\bm{g}\circ(\bm{J}-\bm{X_{1}})\circ\cdots\circ(\bm{J}-\bm{X_{p}})\ \left(\bm{\Xi}^{(t+1)}\right)\right)\oslash$ (23)
$\left(\left(\bm{\Xi}^{(t+1)}\right)^{\top}(\bm{J}-\bm{X_{1}})\circ\cdots\circ(\bm{J}-\bm{X_{p}})\left(\bm{\Xi}^{(t+1)}\right)\right),$
$\bm{\Pi_{0}}\left(\bm{\Xi}^{(t)}\right)=\log\left(1-\bm{\pi}^{(t+1)}(1,\bm{0})\right)$
and calculate $\Omega_{ik}^{(t)}\left(\bm{0},\bm{0},\bm{\Xi}^{(t)}\right)$ for $i=1,\ldots,N$ and $k=1,\ldots,K$.
$\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)~{}$ $=~{}\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left(g_{ij}\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{kl}^{(t)}(1,\bm{x}_{ij})}{\pi_{kl}^{(t)}(0,\bm{0})}\right)+g_{ji}\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{lk}^{(t)}(1,\bm{x}_{ji})}{\pi_{lk}^{(t)}(0,\bm{0})}\right)\right)$
$+\,\sum_{j\neq i}^{N}\sum_{l=1}^{K}\left((1-g_{ij})\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{kl}^{(t)}(0,\bm{x}_{ij})}{\pi_{kl}^{(t)}(0,\bm{0})}\right)+(1-g_{ji})\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{lk}^{(t)}(0,\bm{x}_{ji})}{\pi_{lk}^{(t)}(0,\bm{0})}\right)\right)$
$=\sum_{\bm{x}\in\mathcal{X}}\sum_{j\neq i}^{N}\sum_{l=1}^{K}g_{ij}\,\mathbb{I}(\bm{x}_{ij}=\bm{x})\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{kl}^{(t)}(1,\bm{x})}{\pi_{kl}^{(t)}(0,\bm{0})}\right)$
$+\,g_{ji}\,\mathbb{I}(\bm{x}_{ji}=\bm{x})\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{lk}^{(t)}(1,\bm{x})}{\pi_{lk}^{(t)}(0,\bm{0})}\right)+(1-g_{ij})\,\mathbb{I}(\bm{x}_{ij}=\bm{x})\,\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{kl}^{(t)}(0,\bm{x})}{\pi_{kl}^{(t)}(0,\bm{0})}\right)$
+ $+\,1-g_{ji})\,\mathbb{I}(\bm{x}_{ji}=\bm{x})\,\xi_{jl}^{(t)}\left(\log\,\dfrac{\pi_{lk}^{(t)}(0,\bm{x})}{\pi_{lk}^{(t)}(0,\bm{0})}\right)$
@@
$\bm{\Pi}\left(d,\bm{x},\bm{\Xi}^{(t)}\right)\coloneqq\begin{bmatrix}\log\frac{\pi_{11}^{(t)}(d,\bm{x})}{\pi_{11}^{(t)}(d,\bm{0})}&\log\frac{\pi_{12}^{(t)}(d,\bm{x})}{\pi_{12}^{(t)}(d,\bm{0})}&\dots&\log\frac{\pi_{1K}^{(t)}(d,\bm{x})}{\pi_{1K}^{(t)}(d,\bm{0})}\\ \log\frac{\pi_{21}^{(t)}(d,\bm{x})}{\pi_{21}^{(t)}(d,\bm{0})}&\log\frac{\pi_{22}^{(t)}(d,\bm{x})}{\pi_{22}^{(t)}(d,\bm{0})}&\dots&\log\frac{\pi_{2K}^{(t)}(d,\bm{x})}{\pi_{2K}^{(t)}(d,\bm{0})}\\ \vdots&\vdots&\ddots&\vdots\\ \log\frac{\pi_{K1}^{(t)}(d,\bm{x})}{\pi_{K1}^{(t)}(d,\bm{0})}&\log\frac{\pi_{K2}^{(t)}(d,\bm{x})}{\pi_{K2}^{(t)}(d,\bm{0})}&\dots&\log\frac{\pi_{KK}^{(t)}(d,\bm{x})}{\pi_{KK}^{(t)}(d,\bm{0})}\end{bmatrix},$
$\bm{\Gamma}(d)\coloneqq\begin{cases}\bm{g}&(d=1)\\ \bm{J}-\bm{g}&(d=0)\end{cases},$